The cheetah can maintain its maximum speed
for only 7.5 s.
What is the minimum distance the gazelle
must be ahead of the cheetah to have a chance
of escape? (After 7.5 s the speed of cheetah is
less than that of the gazelle.)
Answer in units of m.
1
Expert's answer
2019-10-07T10:41:09-0400

A cheetah can run at a maximum speed 108 km/h and a gazelle can run at a maximum

speed of 75.9 km/h.


The velocity of the cheetah in meters per second is:

"V_{oc}=108\frac{km}{h}*\frac{1000m}{1km}*\frac{1h}{3600s}\\ 
V_{oc}=30\frac{m}{s}"


The velocity of the gazelle in meters per second.

"V_{og}=75.5\frac{km}{h}*\frac{1000m}{1km}*\frac{1h}{3600s}\\ 
V_{og}=21.1\frac{m}{s}"


Establishing a reference system where initial the cheetah.


The position of the cheetah is given by:


"X_{C}=X_{oc}+V_{oc}*t"


Where:

  • Initial position "X_{0}=0m"
  • Time "t=7.5s"
  • velocity "V_{oc}=30\frac{m}{s}"


Expression for the position."X_{C}=V_{oc}*t"


The position for the gazelle is given by:


"X_{g}=X_{og}+V_{og}*t"


Where:

  • Initial position "X_{og}=??"
  • Velocity of the gazelle "V_{og}=21.1\frac{m}{s}"
  • Time "t=7.5s"


Expression."X_{g}=X_{og}+V_{og}*t"


When the cheetah barely reaches the gazelle it is fulfilled that:


"X_{c}=X_{g}"


replacing expressions


"V_{oc}*t=X_{og}+V_{og}*t"


resolving for the initial position of the gazelle


"V_{oc}*t=X_{og}+V_{og}*t \\ V_{oc}*t=X_{og}+V_{og}*t \\ X_{og} = V_{oc}*t-V_{og}*t \\ X_{og} = 30\frac{m}{s}*7.5s-21.1 \frac{m}{s}*7.5s \text{ Numerically evaluating }\\X_{og}=66.75m"


the minimum distance the gazelle must be ahead of the cheetah to have a chance of escape:

"\boxed{X_{og}=66.75m}"